OpenStudy (anonymous):

what are the coordinates of the foci of the conic section shown below ((x+2)^2/16 )-((y-3)^2/9)=1

OpenStudy (anonymous):

this is a hyperbola.. center at (-2, 3) c^2 = a^2 + b^2 where a=4, b=3 so c=5. your foci is c=5 units above and below the center of the hyperbola.

OpenStudy (anonymous):

1) Given equation is that of a hyperbola with vertical transverse axis of the standard form: (y-k)^2/a^2-(x-h)^2/b^2=1, (h,k) being the (x,y) coordinates of the center. 2) For given equation: center: (-2,3) Vertices: a^2=9 a=3 length of transverse axis=2a=6 vertices: (-2,3±a)=(-2,3±3)=(-2,6) and (-2,0) 3) Foci: b^2=16 b=4 c^2=a^2+b^2=9+16=25 c=5

OpenStudy (anonymous):

See the graph. The axis containing the vertices and the Foci is parallel to the x-axis.

OpenStudy (anonymous):

(-6,3) and (2,3)

OpenStudy (anonymous):

(-6,3) and (2,3) are the vertices.

OpenStudy (anonymous):

hmm.. my bad... thanks @eliassaab :) your foci should be 5 units left/right of the center (-2, 3)